# SOOD++: Leveraging Unlabeled Data to Boost Oriented Object Detection

### v1

@@ 6.1.3 Results on the HRSC2016 dataset
This part evaluates the proposed method on another representative aerial scenes dataset, i.e., HRSC2016 [41]. Notably, the commonly utilized mAP$(07)$ metric fails to adequately show the performance discrepancies among different methodologies. Consequently, we use more stringent evaluation metrics to provide a more precise performance evaluation, including AP85 and mAP50:90.
@@
Fully labeled data. Similar to the above datasets, we also compare the performance under the fully labeled data setting. As shown in Tab. V, our method outperforms previous SSOD methods [35, 76] and surpasses our supervised baseline by a large margin with 8.90 AP85.
Figure 8: Some typical examples from the DOTA-V1.5 dataset. From left to right, there are ground truth, supervised baseline (rotated-FCOS[56]), the SOTA method ARSL [35], and our proposed SOOD++. Row 2 refers to scenarios with extreme aspect ratios case. Row 3 and row 4 are two extremely dense scenes. The red dashed and solid red circles represent false negative and false positive, respectively.
TABLE V: Experimental results on the HRSC2016 (val set) dataset with single-scale training and testing under fully labeled data. Numbers before the arrow are the supervised baseline (rotated-FCOS).
Method Publication AP85
Dense Teacher [76] ECCV 2022 49.50 $\xrightarrow{+3.20}$ 52.70
ARSL [35] CVPR 2023 49.50 $\xrightarrow{+5.00}$ 54.50
SOOD++ (ours) - 49.50 $\xrightarrow{+8.90}$ 58.40
TABLE VI: The generalization ability of our SOOD++. The results are reported with single-scale training and testing.
@@
Oriented R-CNN [62] ICCV 2021 Supervised 67.26
@@ 6.2 Visualization analysis
We further visualize the qualitative results in Fig. 8. These samples are selected from typical scenes on the DOTA-V1.5 dataset (validation set). Specifically, the first sample is a sparse and relatively simpler case, and most semi-supervised methods can work well. However, when facing challenging cases, e.g., the second sample is the large aspect ratio case, and the third and fourth samples are two dense and small cases, we can find that compared with the SOTA method [35], our method presents better visualization. We believe this is because SOOD++ can exploit more potential semantic information from unlabeled data by leveraging the geometry prior and global consistency. These qualitative results further highlight the effectiveness of our method.
6.3 The generalization of our method
This section explores the generalization capabilities of our method from two perspectives: 1) We evaluate our method on different detectors, e.g., CFA [19] and Oriented R-CNN [62]; 2) We conduct experiments on the semi-supervised multi-view oriented 3D object detection to demonstrate the generalization on 3D scenes.
@@ 6.3.1 The generalization on different detectors
To examine the adaptability of our method, we conduct experiments on a suite of representative oriented object detectors under the fully labeled data setting. Specifically, we choose the CFA [19] and Oriented R-CNN [62] for evaluation, as shown in Tab. VI(a). When trained with the full DOTA-V1.5 training data, the fully supervised Oriented R-CNN [62] achieves 67.26 mAP, which is indeed a superior performance. Even so, our method improves it by 2.24 mAP and reaches 69.50 mAP by leveraging the unlabeled data. In addition, our method also reports 2.74 mAP improvement on the CFA [19]. These results are significant, demonstrating our method’s ability to seamlessly integrate with diverse detectors and effectively leverage the intrinsic strengths of the proposed key designs.
We further implement a strong supervised baseline (oriented R-CNN with EVA pre-trained weight [12]) and combine it with our SOOD++ to compare with SOTA oriented detectors [32, 63, 3], as shown in Tab. VI(b). Despite the strong baseline achieving 70.66 mAP on the DOTA-V1.5 test set, our approach shows a 1.82 mAP improvement by leveraging unlabeled data [^3] We utilize DOTA-V2.0 [11] dataset as the unlabeled data in this setting.. It outperforms the PKINet [3] by a notable margin, pushing the new state-of-the-art.
TABLE VII: The generalization on the oriented 3D object detection dataset (nuScenes [2] val set). All semi-supervised methods use the same detector (BEVDet [24]).
@@
TABLE VIII: The effectiveness of our key components: Simple Instance-aware Dense Sampling (SIDS) strategy, Geometry-aware Adaptive Weighting (GAW) loss, and Noise-driven Global Consistency (NGC).
Sampling strategy GAW NGC mAP
10% 20% 30%
From Dense Teacher [76] - - 46.90 53.93 57.86
From SOOD [23] - - 47.24 54.07 57.74
SIDS - - 48.21 55.54 58.69
SIDS - 49.27 56.45 60.11
- 49.32 56.63 59.76
50.48 57.44 61.51
@@
TABLE IX: The analysis of our Simple Instance-aware Dense Sampling (SIDS) strategy. The experiments are equipped with GAW and NGC.
@@
Figure 9: Qualitative results of our method on semi-supervised multi-view oriented 3D object detection task. The above multi-view images show the 3D detection results (green boxes). We also draw the predicted results on the BEV view (the lower part) for better visualizations.
We find that the multi-view oriented 3D objects are very similar to the objects from aerial scenes, as both of them have arbitrary orientations and are usually dense. In addition, the advanced multi-view oriented 3D object detection methods usually detect 3D objects by transforming the perspective from the camera view to the Bird’s Eye View (BEV), similar to aerial scenes captured from a BEV. Thus, to further demonstrate the generalization of our method, we conduct experiments on the most convincing large-scale multi-view oriented 3D object detection dataset, i.e., nuScenes [2]. As shown in Tab. VII, our method reports superior performance, consistently outperforming the supervised baseline and ARSL [35] by a large margin in terms of mAP and NDS metrics under both partially and fully labeled data settings.
In addition, we provide some visualization to present the effectiveness of our method, as shown in Fig. 9. These quantitative and qualitative results prove SOOD++’s ability to be generalized to semi-supervised multi-view oriented 3D object detection.
6.4 Analysis and Ablation Study
The following experiments are conducted on the large-scale DOTA-V1.5 dataset (val set). Unless specified, all the ablation experiments are performed using 10% of labeled data with single-scale training and testing.
6.4.1 The effectiveness of key components
We first study the effects of the proposed key components, including the Simple Instance-aware Dense Sampling (SIDS) strategy, Geometry-aware Adaptive Weighting (GAW) loss, and Noise-driven Global Consistency (NGC), the results listed in Tab. VIII.
There are various dense sampling strategies for constructing dense pseudo-labels. Dense Teacher [76] and our conference version [23] sample dense pseudo-labels from the entire predicted score maps and box area of the foreground, respectively. This work not only considers the easy pseudo-labels of the foreground but also mines the potential pseudo-labels of the background, leading to the proposal of SIDS. Although our SIDS is relatively simple, compared with the strategies in [76, 23], we demonstrate notable improvement, which indicates the constructed higher-quality pseudo-labels of our method.
Based on SIDS, when GAW or NGC is individually incorporated, an increase in mAP is observed. It demonstrates the effectiveness of constructing the one-to-one or many-to-many (global) consistency. Finally, when combining all our key components (NGC, GAW, and SIDS), the improvement will be amplified, achieving satisfactory 50.48, 57.44, and 61.51 mAP under the 10%, 20%, and 30% labeled data settings, respectively. It is reasonable, as the one-to-one and many-to-many relationships are highly complementary. These detailed ablation studies prove the consistent improvements and the value of each component.
@@ 6.4.2 The analysis of Simple Instance-aware Dense Sampling (SIDS) strategy
In this part, we discuss the design of our Simple Instance-aware Dense Sampling (SIDS) strategy, which is used to construct high-quality pseudo-dense labels.
The effect of sample ratio $\mathcal{R}$. Tab. IX(a) studies the effect of the sample ratio $\mathcal{R}$ used in the foreground. The larger the number, the more pseudo-labels are sampled, and vice versa. We can see that the best performance, 50.48 mAP, is achieved when the sample ratio $\mathcal{R}$ is set to 0.25. We hypothesize that this value ensures a good balance between positive and negative pseudo-labels. Increasing it will introduce more noise pseudo-labels that harm the training process, and decreasing it will lead to the loss of some positive pseudo-labels and failure in learning. When set to 1.0, all pixel-level predictions within the instance box are sampled, which inevitably introduces a lot of noise dense pseudo-labels, leading to poor performance.
The filtering criteria used in the background. As discussed in Sec. 4.2, there are many potential hard objects that merge with the background, which are hard to extract as pseudo-labels. Thus, to construct comprehensive dense pseudo-labels, we utilize the predicted IoU score as an auxiliary metric to help sample dense pseudo-labels from the background. We study the effect of using different auxiliary metrics to mine the potential labels, including classification and centeredness scores of default rotated-FCOS [56], as shown in Tab. IX(b). The results prove the necessity of using IoU as the filtering criteria for mining the hard case of background. We argue the main reason is that the IoU score directly reflects the alignment between the predicted and ground truth, making it a more comprehensive metric than the other two.
The effect of threshold $\mathcal{T}_{h}$. We also discuss the effect of the IoU threshold $\mathcal{T}_{h}$, as shown in Tab. IX(c). It is insensitive to the chosen value, and we empirically find that the threshold set as 0.1 can achieve a better performance.
TABLE X: The analysis of our Geometry-aware Adaptive Weighting (GAW) loss. The experiments are equipped with SIDS and NGC.
@@
(b) The effect of $\psi$.
$\psi$ mAP
w/o GAW 49.32
1 49.83
10 49.95
50 50.48
100 50.13
TABLE XI: The analysis of our Noise-driven Global Consistency (NGC). The experiments are equipped with SIDS strategy and GAW loss.
(a) The effect of different global alignments.
@@
(c) The effect of noisy intensity $\beta$.
@@ 6.4.3 The analysis of Geometry-aware Adaptive Weighting (GAW) loss
The effect of geometry information in the GAW loss. In GAW, we utilize the intrinsic geometry information (e.g., orientation gap and aspect ratios) to construct the geometry-aware modulating factor (Eq. 5). To investigate their influence, we conduct experiments on different combinations of the GAW loss. As shown in Tab. X(a), without the GAW, we just achieve 49.32 mAP. By individually utilizing the orientation gap or aspect ratio to construct the GAW, they report similar performance gains. When we combine the two information, the 50.48 mAP performance is reached. This substantial improvement confirms that using both geometric information together provides a more thorough understanding of the oriented objects’ characteristics, leading to better detection performance.
The effect of hyper-parameter $\psi$. We also study the influence of the hyper-parameter $\psi$ in GAW. As shown in Tab. X(b), we set $\psi$ to 1 and get the performance of 49.83 mAP. As $\psi$ increases, the performance of our method improves when $\psi$ varies from 1 to 50. However, further increasing it to 100 slightly hurt the performance. Therefore, we set it to 50 by default. For this observation, we conjecture that increasing the weight $\psi$ will enlarge the influence of geometry information but also amplify the impact of the teacher’s inaccurate labels. Notably, GAW consistently provides performance gains regardless of the $\psi$ value.
@@
The effect of different global alignments. As mentioned in Sec. 4.4, we treat the output of teacher and student as two global distributions and propose to use random noise to disturb the distributions. Then, we make alignment from multi-perspective, including the alignment ($\mathcal{L}_{GC}(\mathbf{d}^{t},\mathbf{d}^{s})$) between the original teacher and the original student, the alignment ($\mathcal{L}_{GC}(\widetilde{\mathbf{d}}^{t},\widetilde{\mathbf{d}}^{s})$) between the disturbed teacher and the disturbed student, and the alignment ($\mathcal{L}_{plan}$) between two OT plans generated by the two former alignments. Thus, we first study the effectiveness of using different alignments. As shown in Tab. XI(a), only using the $\mathcal{L}_{GC}(\mathbf{d}^{t},\mathbf{d}^{s})$, we achieve 49.83 mAP, 0.56 mAP gain compared with the method that equipped with SIDS strategy and GAW loss, indicating the necessary of many-to-many relationship. As expected, by further aligning the two disturbed distributions ($\mathcal{L}_{GC}(\widetilde{\mathbf{d}}^{t},\widetilde{\mathbf{d}}^{s})$) and two transport plans ($\mathcal{L}_{plan}$), we obtain a noticeable improvement, finally resulting in 50.48 mAP. We argue that using random noise to disturb the global distribution mitigates the detrimental effects of label noise and reinforces the model’s ability to learn robust features from the unlabeled data.
@@
The effect of different cost maps. To construct the cost map for solving the OT problem, we consider the spatial distance and the score difference of each possible matching pair. Thus, this part studies the effects of different cost maps of optimal transport in NGC. As shown in Tab. XI(b), individually utilizing spatial distance or score difference results in a maximum improvement of 0.82 mAP, indicating that relying on a single type of information is insufficient for capturing the comprehensive global prior necessary for effective learning. When both spatial distance and score difference are integrated into the cost maps, the improvement is amplified to 1.21 mAP. It highlights the mutually beneficial relationship between these two factors. With their help, we effectively model the many-to-many relationship between the teacher and the student, providing informative guidance to the model.
The effect of $\beta$ and $\mathcal{T}_{g}$. The former hyper-parameter $\beta$ controls the intensity of adding noise, and the latter $\mathcal{T}_{g}$ determines how long a sequence can be considered as a global distribution. As listed in Tab. XI(c) and Tab. XI(d), we empirically find that the best performance is achieved when the values of $\beta$ and $\mathcal{T}_{g}$ are set as 0.3 and 150, respectively.
@@
6.5 Limitation and Related Discussion
This paper provides a simple yet solid framework for semi-supervised oriented object detection. Although our method achieves promising results, the usage of aerial objects’ characteristics is limited. Apart from orientation and global layout, many other properties of aerial objects should be considered, e.g., scale variations. In addition, this paper has explored the generalization of semi-supervised oriented 3D object detection tasks. However, some other oriented detection tasks, such as scene text detection, still need to be studied, and we leave these for the future work.

### v2

@@ 5.3.4 Results on the HRSC2016 dataset
This part evaluates the proposed method on another representative aerial scenes dataset, i.e., HRSC2016 [72]. Notably, the commonly utilized mAP(07)(07) metric fails to adequately show the performance discrepancies among different methodologies. Consequently, we use more strict metrics to provide a more precise performance evaluation, including AP85 and mAP50:95.
@@
Fully labeled data. As shown in Tab. 9, SOOD++ continues to demonstrate strong performance even when all annotations are available. Notably, it surpasses the supervised baseline by a substantial margin of 8.9 AP85—far exceeding the improvements brought by prior SSOD methods, which range from only 3.2 to 5.4. This suggests that SOOD++ is not only effective in low-label regimes but also complements full supervision with more accurate and denser pseudo-labels, leading to superior final performance.
5.3.5 Results on the DIOR-R Dataset
We finally evaluate SOOD++ on the DIOR-R dataset [43], a large-scale benchmark with complex aerial scenes.
Partially labeled data. Tab. 12 shows that SOOD++ consistently outperforms prior semi-supervised methods [7, 6] across 10%, 20%, and 30% labeled settings. Notably, our method achieves 63.1 mAP under the 30% split, significantly surpassing Sparse Semi-DETR [6] by 3.3 points and demonstrating superior effectiveness in low-label regimes.
Fully labeled data. Similar to the above datasets, we also compare the performance under the fully labeled data setting. As summarized in Tab. 12, our method achieves 69.7 mAP, representing a 4.0-point gain over the supervised baseline. This indicates the scalability and transferability of SOOD++ across datasets and annotation densities.
TABLE V: The generalization ability of our SOOD++. The results are reported with single-scale training and testing.
@@ 5.3.5 Results on the DIOR-R Dataset
Oriented R-CNN [46] ICCV 21 Supervised 67.26
@@ 5.4 Visualization Analysis
We further visualize the qualitative results in Fig. 10. We choose the dense pseudo-labeling (DPL)-based methods for visualization. Specifically, the first sample is a sparse and relatively simpler case, and most semi-supervised methods can work well. However, when facing challenging cases, e.g., the second sample is the large aspect ratio case, and the third and fourth samples are two dense and small cases, we can find that compared with ARSL [5], our method presents better visualization. We believe this is because SOOD++ can exploit more potential semantic information from unlabeled data by leveraging the geometry prior and global consistency. These qualitative results further highlight the effectiveness of our method.
5.5 The Generalization Ability
This section evaluates the generalization of our method across detectors (e.g., CFA [79], Oriented R-CNN [46]) and a 3D related task (semi-supervised multi-view 3D oriented object detection).
@@ 5.5.1 The generalization on different detectors
To examine the adaptability of our method, we conduct experiments on a suite of representative oriented object detectors under the fully labeled data setting. Specifically, we choose the CFA [79] and Oriented R-CNN [46] for evaluation, as shown in Tab. V(a). When trained with the full DOTA-V1.5 training data, the fully supervised Oriented R-CNN [46] achieves 67.26 mAP, which is indeed a superior performance. Even so, our method improves it by 2.24 mAP and reaches 69.50 mAP by leveraging the unlabeled data. In addition, our method also reports 2.74 mAP improvement on the CFA [79]. These results are significant, demonstrating our method’s ability to seamlessly integrate with diverse detectors and effectively leverage the intrinsic strengths of the proposed key designs.
We further implement a strong supervised baseline (oriented R-CNN with EVA pre-trained weight [82]) and combine it with our SOOD++ to compare with SOTA oriented detectors [36, 80, 81], as shown in Tab. V(b). Despite the strong baseline achieving 70.66 mAP on the DOTA-V1.5 test set, our approach shows a 1.82 mAP improvement by leveraging unlabeled data [^5] We utilize DOTA-V2.0 [71] dataset as the unlabeled data in this setting.. It outperforms the PKINet [81] by a notable margin, pushing the new state-of-the-art.
Figure 13: Qualitative results for semi-supervised multi-view oriented 3D detection. Upper: Multi-view images with predicted 3D boxes. Lower: Corresponding BEV visualization.
We find that the multi-view oriented 3D objects are very similar to the objects from aerial scenes, as both of them have arbitrary orientations and are usually dense. In addition, the advanced multi-view oriented 3D object detection methods usually detect 3D objects by transforming the perspective from the camera view to the Bird’s Eye View (BEV), similar to aerial scenes captured from a BEV. Thus, to further demonstrate the generalization of our method, we conduct experiments on the most convincing large-scale multi-view oriented 3D object detection dataset, i.e., nuScenes [74]. As shown in Tab. VI, our method reports superior performance, outperforming the supervised baseline and ARSL [5] by a large margin in terms of mAP and NDS metrics under both partially and fully labeled data settings. We also provide detailed visualizations to present the effectiveness of our method, as presented in Fig. 13. These quantitative and qualitative results prove SOOD++’s ability to be generalized to semi-supervised multi-view oriented 3D object detection.
TABLE VI: The generalization on the oriented 3D object detection dataset (nuScenes [74] val set). All semi-supervised methods use the same detector (BEVDet [77]).
@@ 5.5.2 The generalization on semi-supervised multi-view oriented 3D object detection
TABLE VII: Performance comparison using different pseudo-labels on the DOTA-V2.0 (val set) and DOTA-V1.5 (val set) with single-scale training and testing.
Source of Pseudo-labels DOTA-V2.0 DOTA-V1.5
10% 20% 30% 10% 20% 30%
Supervised baseline 38.78 46.74 49.62 42.78 50.11 54.79
+ Gemini 2.5 Flash 39.44 47.09 50.17 43.62 50.87 56.03
+ GPT-4.1 41.07 47.95 52.03 44.78 51.59 56.90
+ SOOD++ (ours) 49.14 54.27 56.90 50.48 57.44 61.51
5.6 Discussion on Vision-Language Models
Recent multimodal large language models, such as GPT-4.1 and Gemini 2.5, have inspired interest in their use for pseudo-label generation in visual tasks. To assess their effectiveness for semi-supervised oriented object detection, we generate pseudo-labels using these MLLMs and compare their performance with our SOOD++ framework. As shown in Tab. VII, although MLLMs offer slight gains over the supervised baseline (e.g., +0.66 with Gemini 2.5 and +2.29 with GPT-4.1 under the 10% setting on DOTA-V2.0), their performance remains substantially behind SOOD++, which achieves a remarkable improvement of +10.36 mAP. This significant gap highlights the limitations of current MLLMs in handling spatially complex and densely annotated tasks such as oriented object detection. Note that such similar findings are also observed in general vision tasks, as discussed in a recent work [83], where MLLMs struggle with fine-grained localization. These observations suggest that despite their broad visual-textual understanding, MLLMs lack the spatial precision, orientation sensitivity, and dense region awareness required by this task. In contrast, SOOD++ incorporates task-specific strategies that are explicitly tailored for the challenges of oriented object detection. As a result, it produces substantially more reliable pseudo-labels and delivers superior semi-supervised performance across various data settings.
@@ 5.6 Discussion on Vision-Language Models
TABLE VIII: The effectiveness of Simple Instance-aware Dense Sampling (SIDS) strategy, Geometry-aware Adaptive Weighting (GAW) loss, and Noise-driven Global Consistency (NGC) on DOTA-V2.0 and DOTA-V1.5 (val set).
Sampling strategy GAW NGC DOTA-V2.0 DOTA-V1.5
10% 20% 30% 10% 20% 30%
From Dense Teacher [3] - - 42.73 49.51 53.17 46.90 53.93 57.86
From SOOD [53] - - 43.15 50.07 53.54 47.24 54.07 57.74
SIDS - - 44.32 51.48 54.45 48.21 55.54 58.69
SIDS - 47.65 53.22 55.71 49.27 56.45 60.11
- 47.82 53.15 55.98 49.32 56.63 59.76
49.14 54.27 56.90 50.48 57.44 61.51
5.7 Ablation Study
All the ablation experiments are performed using 10% of labeled data with single-scale training and testing.
5.7.1 The effectiveness of key components
We first study the effects of the key components: the Simple Instance-aware Dense Sampling (SIDS) strategy, Geometry-aware Adaptive Weighting (GAW) loss, and Noise-driven Global Consistency (NGC), as shown in Tab. VIII: 1) There are various dense sampling strategies for constructing dense pseudo-labels. Dense Teacher [3] and our conference version [53] sample dense pseudo-labels from the entire predicted score maps and the box area of the foreground, respectively. This work not only considers the easy pseudo-labels of the foreground but also mines the potential pseudo-labels of the background, leading to the proposal of SIDS. Although our SIDS is relatively simple, compared with the strategies in [3, 53], we demonstrate notable improvement, indicating the higher-quality pseudo-labels of our method. 2) Based on SIDS, when GAW or NGC is individually incorporated, an increase in mAP is observed. This demonstrates the effectiveness of constructing the one-to-one or many-to-many (global) consistency. For example, introducing NGC significantly amplifies performance of +3.50, +1.67, and +1.53 mAP on the DOTA-V2.0 dataset, which validates its core principle of reinforcing global representation consistency through multi-perspective alignment. 3) Combining NGC, GAW, and SIDS leads to the best overall performance, with consistent improvements across large-scale DOTA-V2.0 and DOTA-V1.5 datasets: 49.14/50.48, 54.27/57.44, and 56.90/61.51 mAP under 10%, 20%, and 30% label ratios, respectively. It is reasonable, as the one-to-one and many-to-many relationships are highly complementary. These detailed ablation studies prove stable improvements in the value of each component.
Unless specified, the following ablation experiments are conducted on the large-scale DOTA-V1.5 dataset (val set).
5.7.2 The component improvements from SOOD to SOOD++
To intuitively compare the improvement with the conference version [53] (SOOD), we choose SOOD as the baseline and replace each component one at a time with its corresponding new counterpart, as shown in Tab. IX. We can observe that every new refinement module plays a crucial role in the performance boost. For example, replacing the previous sampling strategy with SIDS results in a significant improvement of +0.81 mAP at 10%, +0.67 mAP at 20%, and +1.13 mAP at 30%. Similarly, replacing RAW with GAW and GC with NGC also led to notable improvements. When these all-new modules are combined, we achieve significant performance improvements: +1.85, +1.86, and +2.28 under 10%, 20%, and 30% labeled data settings. These results clearly demonstrate the crucial contribution of each improvement in SOOD++ and how they work together to provide a more effective solution.
TABLE IX: Component-wise comparison between SOOD [53] and SOOD++ on DOTA-V1.5 (val set).
Method 10% 20% 30%
SOOD [53] 48.63 55.58 59.23
Replacing sampling strategy with SIDS 49.44(+0.81) 56.25(+0.67) 60.36(+1.13)
Replacing RAW with GAW 49.38(+0.75) 56.29(+0.71) 60.03(+0.80)
Replacing GC with NGC 49.61(+0.98) 56.47(+0.89) 60.25(+1.02)
SOOD++ (ours) 50.48(+1.85) 57.44(+1.86) 61.51(+2.28)
@@ 5.7.2 The component improvements from SOOD to SOOD++
TABLE X: The analysis of our Simple Instance-aware Dense Sampling (SIDS) strategy. The experiments are equipped with GAW and NGC.
@@ 5.7.3 The analysis of Simple Instance-aware Dense Sampling (SIDS) strategy
In this part, we discuss the design of our SIDS strategy, which is used to construct high-quality pseudo-dense labels.
The effect of sample ratio $\mathcal{R}$. Tab. X(a) studies the effect of the sample ratio $\mathcal{R}$ used in the foreground. The larger the number, the more pseudo-labels are sampled, and vice versa. We can see that the best performance, 50.48 mAP, is achieved when the sample ratio $\mathcal{R}$ is set to 0.25. We hypothesize that this value ensures a good balance between positive and negative pseudo-labels. Increasing it will introduce more noise pseudo-labels that harm the training process, and decreasing it will lead to the loss of some positive pseudo-labels and failure in learning. When set to 1.0, all pixel-level predictions within the instance box are sampled, which inevitably introduces a lot of noise dense pseudo-labels, leading to poor performance.
The filtering criteria used in the background. As discussed in Sec. 4.2, there are many potential hard objects that merge with the background, which are hard to extract as pseudo-labels. Thus, to construct comprehensive dense pseudo-labels, we utilize the predicted IoU score as an auxiliary metric to help sample dense pseudo-labels from the background. We study the effect of using different auxiliary metrics to mine the potential labels, including classification and centerness scores of default rotated-FCOS [56], as shown in Tab. X(b). The results prove the necessity of using IoU as the filtering criterion for mining the hard cases of background. We argue that the main reason is that the IoU score directly reflects the alignment between the predicted and ground truth, making it a more comprehensive metric.
The effect of IoU threshold $\mathcal{T}_{h}$. This parameter is introduced to select the pixel-level points with high IoU scores from the background to construct the pseudo labels. Setting its value too high may under-sample, and too low may introduce too much noise. Therefore, we set $\mathcal{T}_{h}$ as 0.1, and it works well, as shown in Tab. X(c).
TABLE XI: The analysis of our Geometry-aware Adaptive Weighting (GAW) loss. The experiments are equipped with SIDS and NGC.
@@
(b) The effect of loss balance weight
$\psi$. $\psi$ mAP w/o GAW 49.32 1 49.83 10 49.95 50 50.48 100 50.13
TABLE XII: The analysis of our Noise-driven Global Consistency (NGC). The experiments are equipped with SIDS strategy and GAW loss. $\mathcal{L}{GC}(\mathbf{d}^{t},\mathbf{d}^{s})$, $\mathcal{L}{GC}(\widetilde{\mathbf{d}}^{t},\widetilde{\mathbf{d}}^{s})$, and $\mathcal{L}_{plan}$ denote alignments between original teacher-student, disturbed teacher-student, and the corresponding optimal transport plans, respectively.
(a) The effect of different alignments.
@@
TABLE XIII: Ablation study on the noisy intensity $\beta$ and global number threshold $\mathcal{T}_{g}$.
(a) The effect of intensity $\beta$.
@@ 5.7.4 The analysis of Geometry-aware Adaptive Weighting (GAW) loss
The effect of geometry information in the GAW loss. In GAW, we utilize the intrinsic geometry information (e.g., orientation gap and aspect ratios) to construct the geometry-aware modulating factor (Eq. 5). To investigate their influence, we conduct experiments on different combinations of the GAW loss. As shown in Tab. XI(a), without the GAW, we just achieve 49.32 mAP. By individually utilizing the orientation gap or aspect ratio to construct the GAW, they report similar performance gains. When we combine the two types of information, the 50.48 mAP performance is reached. This substantial improvement confirms that using both geometric information together provides a more thorough understanding of the oriented objects’ characteristics, leading to better detection performance.
The effect of loss balance weight $\psi$. We also study the influence of the hyper-parameter $\psi$ in GAW. As shown in Tab. XI(b), we set $\psi$ to 1 and get the performance of 49.83 mAP. As $\psi$ increases, the performance of our method improves when $\psi$ varies from 1 to 50. However, further increasing it to 100 slightly hurt the performance. Therefore, we set it to 50 by default. For this observation, we conjecture that increasing the weight $\psi$ will enlarge the influence of geometry information but also amplify the impact of the teacher’s inaccurate labels. Notably, GAW consistently provides performance gains regardless of the $\psi$ value.
@@
The effect of different global alignments. As mentioned in Sec. 4.4, we treat the output of the teacher and student as two global distributions and propose to use random noise to disturb the distributions. Then, we make alignment from multi-perspective, including the alignment ($\mathcal{L}_{GC}(\mathbf{d}^{t},\mathbf{d}^{s})$) between the original teacher and the original student, the alignment ($\mathcal{L}_{GC}(\widetilde{\mathbf{d}}^{t},\widetilde{\mathbf{d}}^{s})$) between the disturbed teacher and the disturbed student, and the alignment ($\mathcal{L}_{plan}$) between two OT plans generated by the two former alignments. Thus, we first study the effectiveness of using different alignments. As shown in Tab. XII(a), only using the $\mathcal{L}_{GC}(\mathbf{d}^{t},\mathbf{d}^{s})$, we achieve 49.83 mAP, 0.56 mAP gain compared with the method that equipped with SIDS strategy and GAW loss, indicating the necessity of a many-to-many relationship. As expected, by further aligning the two disturbed distributions ($\mathcal{L}_{GC}(\widetilde{\mathbf{d}}^{t},\widetilde{\mathbf{d}}^{s})$) and two transport plans ($\mathcal{L}_{plan}$), we obtain a noticeable improvement, finally resulting in 50.48 mAP. We argue that using random noise to disturb the global distribution mitigates the detrimental effects of label noise and reinforces the model’s ability to learn robust features from the unlabeled data.
@@
The effect of different cost maps. To construct the cost map for solving the OT problem, we consider the spatial distance and the score difference of each possible matching pair. Thus, this part studies the effects of different cost maps of optimal transport in NGC. As shown in Tab. XII(b), individually utilizing spatial distance or score difference results in a maximum improvement of 0.82 mAP, indicating that relying on a single type of information is insufficient for capturing the comprehensive global prior necessary for effective learning. When both spatial distance and score difference are integrated into the cost maps, the improvement is amplified to 1.21 mAP. It highlights the mutually beneficial relationship between these two factors. With their help, we effectively model the many-to-many relationship between the teacher and the student, providing informative guidance to the model.
The effect of hyper-parameters $\beta$ and $\mathcal{T}_{g}$. We conduct ablation studies to examine two hyper-parameters in our proposed NGC module: the noisy intensity $\beta$ and the global number threshold $\mathcal{T}_{g}$. As shown in Tab.XIII(a), if $\beta$ is set too low (e.g., 0.1), the injected noise is insufficient to effectively encourage the model to escape local minima, limiting generalization. Conversely, excessively high values (e.g., 0.5) can disrupt global consistency, reducing performance. Therefore, we choose an intermediate value $\beta=0.3$ to strike a balance. Regarding the threshold $\mathcal{T}_{g}$ (Tab.XIII(b)), lower values (e.g., 100) incorrectly treat limited sampled points as global distributions, insufficiently capturing global patterns, while higher values (e.g., 250) overly restrict scenarios benefiting from global consistency. Thus, we select $\mathcal{T}_{g}=150$.